OpenStudy (anonymous):

[8.02] Factor completely: ax + 2ay + bx + 2by (x + b)(a + 2y) (x + a)(x + b) (x + 2y)(a + b) Prime

OpenStudy (anonymous):

could you giev me a straight answer please?

OpenStudy (blurbendy):

(x + b)(a + 2y) = ax + 2xy (wrong already) (x + a)(x + b) = x^2 (wrong already) (x + 2y)(a + b) = ax + bx + 2ay + 2by (correct)